OpenStudy (anonymous):

sqrt7 x sqrt5 multiply and simplify. would it stay at sqrt30?

mathslover (mathslover):

what is 7*5 ?

mathslover (mathslover):

\[\large{\sqrt{7}\times \sqrt{5}=\sqrt{7*5} \ne \sqrt{30}}\]

mathslover (mathslover):

it is : \[\large{\sqrt{35}}\]

OpenStudy (anonymous):

Hint: